# PASCAL: Precise and Efficient ANN- SNN Conversion using Spike Accumulation and Adaptive Layerwise Activation

### v1

@@ 4.1 PASCAL: Accuracy Results
Table 3: Energy comparison between ANN and SNN inferred using PASCAL based on the method proposed by Rathi & Roy. The MAC (AC) energy is estimated as $4.6~{}pJ$ ($0.9~{}pJ$(Horowitz, 2014). The spike rate per layer for various models is shown in Appendix A.9.
@@
Architecture (quantization step) Dataset Normalized #$\#OP_{ANN}(a)$ Normalized #$\#OP_{SNN}(b)$ 1 #$\frac{\#OP_{SNN\_Layer1}}{\#OP_{Total}}(c)$ $\frac{Energy_{ANN}}{Energy_{SNN}}$ $(\frac{a*4.6}{c*4.6+(1-c)*b*0.9})$
VGG-16 (L=$4$) CIFAR-10 $1.0$ $0.66$ $0.005$ (refer Appendix A.5) $7.49$
VGG-16 (L=$4$) CIFAR-100 $1.0$ $0.62$ $0.005$ (refer Appendix A.5) $7.96$
VGG-16 (L=$16$) ImageNet $1.0$ $0.73$ $0.006$ (refer Appendix A.6) $6.76$
ResNet-18 (L=$8$) CIFAR-10 $1.0$ $0.86$ $0.008$ (refer Appendix A.7) $5.72$
ResNet-18 (L=$8$) CIFAR-100 $1.0$ $0.91$ $0.008$ (refer Appendix A.7) $5.42$
@@
4.2 PASCAL: Energy Efficiency Results
The intrinsic efficiency of SNNs can be attributed to their ability to perform sparse accumulates (AC), which is 5$\sim$ 5$\times$ 5 × more energy-efficient than the MAC operations performed by ANNs for 32-bit floating point representation (Horowitz, 2014). We estimate the energy efficiency of SNN over ANN by comparing the number of operations between the source ANN (#$\#OP_{ANN}$) and the converted SNN (#$\#OP_{SNN}$), as proposed by Rathi & Roy and formulated as
#$\#OP_{SNN}=SpikeRate_{l}\times\#OP_{ANN}.$ (7)
$SpikeRate_{l}$ is the total spikes in layer $l$ over all timesteps averaged over the number of neurons in the layer. Table 3 shows that SNNs inferred using PASCAL achieves equivalent accuracy as the source ANN while consuming up to 7×7$\times$7 × lower energy. The energy results account for the fact that the first SNN layer alone performs MAC operations since it receives input image with real-valued pixel intensities. The formulae used to compute the ANN and SNN operations are detailed in Appendix A.4.
@@ 4.2 PASCAL: Energy Efficiency Results
AL Metric Model Dataset $\chi$ L Layerwise L $T_{eff}$ SNN Accuracy
Figure  1(a) VGG-16 CIFAR-10 3 1,2,4 [2,1,2$,1$,1,1,4,4,2,4,4] 2.07 93.73 %
Figure  1(b) ResNet-18 CIFAR-10 2 1,4 [4,4,4,1,4,4,4,1,4,1,1,1,1,1,1,1,4] 2.41 95.34%
Figure  1(c) VGG-16 CIFAR-100 3 1,4,8 [8,4,4,1,4,4,1,1 ,1,1,1,4,8,8,8] 3.87 74.07 %
VGG-16 ImageNet 3 16,8,4 [16,16,16,8,8,8,8,8,4,4,4,4,4,4,4] 7.73 71.188
Table 5: SNN accuracy and inference timesteps comparison with other state-of-the-art works
Comparison Work Dataset Model Value of T Accuracy of Comparison Work(%) Accuracy of PASC(%)
(Hao et al., 2023) CIFAR-100 VGG-16 8 76.52 76.77
(Hao et al., 2023) ImageNet ResNet-34 8 74.17 74.31
(Li et al., 2023a) CIFAR-10 ResNet-18 - 94.27 (T=11.51) 96.51 (T=8)
(Li et al., 2023a) ImageNet VGG-16 - 73.30 (T=49.54) 74.22 (T=16)
(Wu et al., 2024) ImageNet ResNet34 16 71.66 74.31
Table 6: SNN accuracy and inference timesteps comparison with QCFS and SEENN-1
PASC formulation with uniform Quantization Step PASC formulation with Adaptive Layerwise (AL) activation
@@
PASC 4 95.61% SEENN-1 2.01 95.08%
CIFAR-100 QCFS 16 72.80% PASCAL 2.41 95.34%
PASC 4 76.06% CIFAR-100 SEENN-1 6.19 65.48%
ResNet-34 ImageNet QCFS 64 72.35% PASCAL 2.35 76.19 %
@@
SEENN-1 29.53 71.84% PASCAL 3.87 74.07%
PASC 8 74.31%
VGG-16 ImageNet QCFS 1024 74.32%
PASC 16 74.22%
@@ 4.3 Adaptive Layerwise (AL) Activation Results
Figure 1 shows the layerwise AL metric (refer Equation 6) for different models and datasets. We perform clustering with various values of $\chi$, and compute the quantization step per layer. We also evaluate the effective inference timesteps, $T_{eff}$, which is specified by
$T_{eff}=\frac{1}{|\mathbf{L_{tot}}|}\cdot\sum_{n=1}^{|\mathbf{L_{tot}}|}L_{n}$ (8)
where $\mathbf{L_{tot}}$ is the total number of layers. Table 4 shows that adapting the quantization step layerwise leads to competitive accuracy using fewer effective timesteps, thereby improving the temporal efficiency. For instance, ResNet-18 provides 95.34% accuracy on CIFAR-10 using $T_{eff}\approx 2.4$, which is comparable to that (96.51%) achieved using $8$ timesteps uniformly for all the layers.
@@ 4.4 Comparison with the State-of-the-Art
We compare our results (shown in Table 6) with the QCFS framework (Bu et al., 2023) and the SEENN-1 method (Li et al., 2023b). Apart from these baselines, we also compare our work with various other state-of-the-art works 5. Our results show that PASCAL yields SNNs that provide competitive accuracy using fewer timesteps, consistently across models and datasets, compared to the QCFS framework (Bu et al., 2023). We report comparable performance to the SEENN-1 method (Li et al., 2023b) on CIFAR-10. However, we demonstrate significant performance improvements, both in terms of accuracy and effective timesteps, for CIFAR-100 and ImageNet. The efficacy of PASCAL improves with dataset complexity since it relies on mathematically equivalent ANN-SNN conversion.
@@ 5 Conclusion
In this paper, we proposed PASCAL, which is a precise and temporally efficient ANN-SNN conversion methodology. We showed that the proposed methodology enables ANNs to be converted into SNNs with a guarantee of mathematical equivalence, thereby leading to near-zero conversion loss. We formally proved that the converted SNN performs the same computations, unrolled over timesteps, as the source ANN with QCFS activation. We demonstrated the energy efficiency of the proposed approach using an analytical model. In addition, we also proposed an Adaptive Layerwise (AL) activation methodology to further improve the temporal efficiency during SNN inference. We presented a systematic and algorithmic approach to choose an optimal precision for each layer, which reduces the effective number of inference timesteps while providing competitive accuracy.
@@ A.1 Proof for equivalence between the PASC formulation and an ANN with QCFS activation
For the input layer, the MatMul Operation performed in the PASC formulation is the same as the corresponding ANN with QCFS activation. Let the output after the MatMul Operation be $X$. Consider a single element $x\in X$. In the case of ANN, the output after Integrate-and-Fire is
@@
Upon further simplification, we get that
$c=\begin{cases}k&\text{if }0<k<L_{n},\text{ and }\\ &(k-\frac{1}{2})\frac{\theta_{n}}{L_{n}}\leq x<(k+\frac{1}{2})\frac{\theta_{n}}{L_{n}}\\ L_{n}&\text{if }x\geq(L_{n}-\frac{1}{2})\frac{\theta_{n}}{L_{n}}\\ 0&\text{if }x<\frac{\theta_{n}}{2\cdot L_{n}}\\ \end{cases}$ (10)
Algorithm 2 specifies the Integrate-and-Fire algorithm for the input layer. It also first generates $\widehat{h}(X)=z^{l}$ (refer Line 3). Subsequently, it generates the final spike train $s^{l}$. We denote the spike train for a given input $x\in X$ by $s^{l}_{x}$. Note that $z^{l}$ only consists of integral multiples of $\theta^{*}=\frac{\theta_{n}}{L_{n}}$. Therefore, after repeated application of the Heaviside step function on $x$, and stacking the output over time, we get that
@@
Now, $\sum_{i=1}^{L_{n}}s^{l}_{x}(i)=c\cdot\frac{\theta_{n}}{L_{n}}=\widehat{h}(x)$, $\forall x\in X$. Therefore, $\sum_{i=1}^{L_{n}}s^{l}(i)=z^{l}$, proving part (a). Using Equation 11, we can see that the $s_{x}^{l}$ contains values from the set $\{0,\frac{\theta_{n}}{L_{n}}\}$ for all values of $x\in X$. This proves part (b) of the theorem.
Proof for Theorem 3.4. We begin with the remark that $\theta_{n}=\lambda^{l}$ after training is complete. Consider a single element $x\in z^{l}$, and corresponding elements $x_{i}\in z^{l}_{i},\forall i\in\{1,2,\cdots L_{n-1}\}$. In an ANN with QCFS Activation (refer Figure 2),
$\widehat{h}(x)=\begin{cases}k\cdot\frac{\theta_{n}}{L_{n}}&\text{if }0<k<L_{n},\text{ and }\\ &(k-\frac{1}{2})\frac{\theta_{n}}{L_{n}}\leq x<(k+\frac{1}{2})\frac{\theta_{n}}{L_{n}}\\ \theta_{n}&\text{if }x\geq(L_{n}-\frac{1}{2})\frac{\theta_{n}}{L_{n}}\\ 0&\text{if }x<\frac{\theta_{n}}{2\cdot L_{n}}\\ \end{cases}$ (12)
After the first stage of Algorithm 1, let the value of $s$ be $k_{1}\cdot\frac{\theta_{n}}{L_{n}}$. If there was a spike in the last timestep, then we obtain an inequality of the form $\frac{\theta_{n}}{2\cdot L_{n}}+\sum_{i}x_{i}-k_{1}\cdot\frac{\theta_{n}}{L_{n}}\geq\frac{\theta_{n}}{L_{n}}$. Upon simplification, we get that
$\sum_{i=1}^{L_{n-1}}x_{i}\geq(k_{1}+\frac{1}{2})\cdot\frac{\theta_{n}}{L_{n}}$ (13)
On the other hand, if there was no spike in the last timestep, we get that
$\sum_{i=1}^{L_{n-1}}x_{i}<(k_{1}+\frac{1}{2})\cdot\frac{\theta_{n}}{L_{n}}$ (14)
Although we have one-directional inequality in either case, we still need more processing to obtain the second direction of the inequality to choose the correct level. We consider the aforementioned two cases.
Case 1. Let Inequality 13 be the final inequality after stage 1. Inhibitory spikes are not possible in this case. The QCFS function clips the activation if the accumulated input $\sum_{i}x_{i}$ is greater than $(L_{n}-\frac{1}{2})\cdot\frac{\theta_{n}}{L_{n}}$. Therefore, to determine the exact level, we would need to run spike without input at least $T_{1}=L_{n}-1$ times to ensure that the correct level is obtained for all inputs. The bound arises because the minimum value of $k_{1}=0$. At the end of stage 2, let the value of $s$ be $k_{2}\cdot\frac{\theta_{n}}{L_{n}}$. Since an inhibitory spike is not possible, we get that
Also, if $k_{2}<L_{n}$, then we also have the inequality
Otherwise, we only have Inequality 15. In effect,
Case 2. If Inequality 14 is the final inequality, we need inhibitory spikes to determine the correct level. This is different from Case 1, because we have an upper bound on the value of $\sum_{i=1}^{L_{n-1}}x_{i}$. In effect, any level below $k_{1}$ should be possible. Since $k_{1}\leq L_{n-1}-1$, this stage requires $T_{2}=L_{n-1}-1$ timesteps. Here, normal spikes are not possible. Using a similar analysis as case 1, we get
Stage 3 is used to generate a final spike train from the value of $s$ obtained after stage 2. This layer is used to generate a stacked output $\widetilde{h}(X)$. As a result of this computation, we get that $\sum_{j}\widetilde{h}(X)[j]=min(max(s,0),L_{n})$. The final number of spikes parametrized by the value of $s$ at the end of stage 2 is given in Table 7. Using 17, 18, we can conclude that
Table 7: Mapping between the value of $s$ at the end of stage 2, and the final number of spikes at the end of stage 3
@@
$0\geq s\geq-(L_{n}-1)$ 00
Figure 2: A sample QCFS activation layer with $L=4$ and $\theta_{n}=0.25$
Proof for Theorem 3.3 We begin by simplifying the right hand side.
@@
$=\sum_{i=1}^{L}\gamma\cdot\frac{z_{i}^{l}*W}{\sqrt{\sigma^{2}+\epsilon}}+\frac{1}{\sqrt{\sigma^{2}+\epsilon}}\sum_{i=1}^{L}\frac{b-\mu}{L}+\sum_{i=1}^{L}\frac{\beta}{L}$ (21)
$=S_{1}+S_{2}+S_{3}$ (22)
@@
$=\frac{\gamma}{\sqrt{\sigma^{2}+\epsilon}}\cdot z^{l}*W$ (24)
$=\frac{\gamma}{\sqrt{\sigma^{2}+\epsilon}}\cdot(z^{l+1}-\beta)\cdot\frac{\sqrt{\sigma^{2}+\epsilon}}{\gamma}-b+\mu$ (25)
$=z^{l+1}-\beta-b+\mu$ (26)
@@
Consider an input $z$ to models $\mathcal{M}$ and $\mathcal{N}$. We invoke Theorems 3.5, 3.4 and 3.3 to show that the invariant 2 holds after the final input layer. In other words, after the final layer, $z^{l}=\sum_{i=1}^{L}z_{i}^{l}$, where $L$ is the quantization step of the final layer $l$. In an SNN, the output probabilities are derived by taking the mean over time-steps after the final layer. Therefore, the tensor used for classification in the PASC formulation is $z^{*}=\frac{1}{L}\cdot\sum_{i=1}^{L}z_{i}^{l}=\frac{1}{L}\cdot z^{l}$. Therefore,
@@
$=\frac{\frac{z^{l}[c]}{L}}{\frac{\sum_{c}z^{l}[c]}{L}}$ (32)
$=\frac{z^{l}[c]}{\sum_{c}z^{l}[c]}$ (33)
$=f_{\mathcal{N}}^{z}(c)$ (34)

### v2

@@ 4.1 PASCAL: Accuracy Results
4.2 PASCAL: Energy Efficiency Results
4.2.1 Energy of PASCAL-SNN with respect to vanilla SNN
SNNs gain their energy efficiency as a result of replacing compute-intensive Multiply-and-Accumulate (MAC) operations by more energy-efficient Accumulate (AC) operations. The energy consumed by the entire network, $E_{SNN}$, can be split into two parts as follows.
$E_{SNN}=E_{AC}+E_{IF}$ (7)
Unlike previous works (rathi2021diet), we also take into account the energy consumed by the Integrate-and-Fire layer ($E_{IF}$). Similar to Equation 7, we express the total energy consumed by the corresponding SNN trained using the PASCAL approach as shown below.
$E_{PASC\_SNN}=E_{PASC\_AC}+E_{PASC\_IF}$ (8)
We note that the energy consumed by AC operations does not change in PASCAL compared to the vanilla SNN. This is because for a given quantization step $L$, both networks have a MatMul layer that is unrolled across $T=L$ timesteps. Therefore, we have $E_{PASC\_AC}=E_{AC}$. The only difference lies in the IF layer, which is more complex in PASCAL. Consider a network with uniform quantization step $L$. According to Algorithm 1, the number of soft reset operations is 3L13L-1, while the number of spike accumulation operations (i.e., addition and subtraction of the spike counter) is 2L12L-1. In total, each IF layer in PASC framework performs (5L2)(5L-2) operations, yielding $E_{PASC\_IF}\approx(5L-2)\cdot E_{IF}$. We quantify the excess energy consumed by the PASC SNN by comparing it with the energy consumed by the vanilla SNN. Specifically, we define the per-layer energy ratio $r_{E}^{l}$ as
$r^{l}_{E}=\frac{E^{l}_{PASC\_SNN}}{E^{l}_{SNN}}=\frac{E^{l}_{AC}+(5L-2)\cdot E^{l}_{IF}}{E^{l}_{AC}+L\cdot E^{l}_{IF}}.$ (9)
We now analytically compute the ratio from Equation 9 for both convolutional and fully connected layers. To facilitate this, we introduce an auxiliary ratio $r^{\prime}$ that represents the number of operations in an IF layer relative to those in the preceding MatMul layer. This allows us to express $r^{l}_{E}$ in terms of $r^{\prime}$ as
$r^{l}_{E}=\frac{1+(5L-2)\cdot r^{\prime}}{1+L\cdot r^{\prime}}.$ (10)
Finally, we proceed to compute $r^{\prime}$ layerwise separately for both the convolution and fully connected layers, using notation outlined in Table A.8
Convolution Layer
We assume a constant quantization step $L$ for further calculations in this section. Each IF layer performs $L\times C_{0}\times H_{0}\times W_{0}$ accumulation operations. Therefore, we compute the ratio between the number of operations in the IF layer to the number of operations in the previous convolution layer as
$r^{\prime}_{conv}=\frac{E^{l}_{IF}}{E^{l}_{AC}}=\frac{\#OP^{l}_{IF}}{\#OP^{l}_{Conv}}\approx\frac{1}{C_{i}\cdot K_{h}\cdot K_{w}\cdot SpikeRate_{l}},$ (11)
where $SpikeRate_{l}$ is the total spikes in layer $l$ across all timesteps averaged over the number of neurons in the layer.
Fully Connected (FC) Layer
Following a similar reasoning to the previous section, we can deduce $r^{\prime}$ for the fully connected layer as
$r^{\prime}_{FC}=\frac{E^{l}_{IF}}{E^{l}_{AC}}=\frac{\#OP^{l}_{IF}}{\#OP^{l}_{FC}}\approx\frac{1}{C_{i}\cdot SpikeRate_{l}}.$ (12)
We use the above equations defining $r^{\prime}$ and $r_{E}^{l}$ to find the normalized number of inference timesteps $T_{norm}$ in the PASC framework as
$T_{norm}=\left(\frac{1}{|\mathbf{L_{tot}}|}\cdot\sum_{l}r^{l}_{E}\right)\cdot T,$ (13)
where $\mathbf{L_{tot}}$ is the total number of layers and $T$ is the number of time-steps of the corresponding vanilla SNN. For this computation, we assume that the $SpikeRate_{l}=0.75$ for all layers. In addition, we also define the overall energy ratio as
$r_{E}=\frac{E^{tot}_{PASC\_SNN}}{E^{tot}_{SNN}},$ (14)
which characterizes the overall energy cost of our method with respect to the traditional SNN. We show in Appendix A.10 that the ratio $r_{E}$ is as low as 1.001 in practice, indicating that increasing the number of IF operations does not adversely affect the overall energy consumption. This shows the ability of the proposed PASC framework to provide lossless ANN-SNN conversion for negligible increase in energy consumption.
4.2.2 Energy comparison of vanilla SNN with ANN
Having established that PASCAL preserves the energy benefits of vanilla SNNs, we now analytically compute the energy efficiency of an SNN relative to its source ANN. We estimate this energy efficiency by comparing the number of operations between the source ANN ($OP_{ANN}$) and the converted SNN ($OP_{SNN}$), following the approach proposed by rathi2021diet and formulated as
$OP_{SNN}=SpikeRate_{l}\times\#OP_{ANN}.$ (15)
Table 3 shows that converted SNNs are up to $7\times$ more energy efficient than the corresponding source ANNs. The energy results account for the fact that the first SNN layer alone performs MAC operations since it receives input image with real-valued pixel intensities. The formulae used to compute the ANN and SNN operations are detailed in Appendix A.4.
@@ 4.2.2 Energy comparison of vanilla SNN with ANN
Table 3: Energy comparison between ANN and SNN based on the method proposed by rathi2021diet. The FP32 MAC (AC) energy is estimated as 4.6pJ4.6~pJ (0.9pJ0.9~pJ) and the INT8 MAC(AC) energy is estimated as 0.23pJ0.23~pJ (0.03pJ0.03~pJ(horowitz20141). The spike rate per layer for various models is shown in Appendix A.9.
@@
Architecture (quantization step) Dataset Normalized $OP_{ANN}(a)$ Normalized $OP_{SNN}(b)$ 1 $\frac{\#OP_{SNN\_Layer1}}{\#OP_{Total}}(c)$ $\frac{Energy_{ANN}}{Energy_{SNN}}(FP32)$ $(\frac{a*4.6}{c*4.6+(1-c)*b*0.9})$ $\frac{Energy_{ANN}}{Energy_{SNN}}(INT8)$ $(\frac{a*0.23}{c*0.23+(1-c)*b*0.03})$
VGG-16 (L=$4$) CIFAR-10 $1.0$ $0.66$ $0.005$ (Appendix A.5) $7.49$ 11.03
VGG-16 (L=$4$) CIFAR-100 $1.0$ $0.62$ $0.005$ (Appendix A.5) $7.96$ 11.70
VGG-16 (L=$16$) ImageNet $1.0$ $0.73$ $0.006$ (Appendix A.6) $6.76$ 9.94
ResNet-18 (L=$8$) CIFAR-10 $1.0$ $0.86$ $0.008$ (Appendix A.7) $5.72$ 8.38
ResNet-18 (L=$8$) CIFAR-100 $1.0$ $0.91$ $0.008$ (Appendix A.7) $5.42$ 7.95
@@
AL Metric Model Dataset $\chi$ L Layerwise L $T_{eff}$ Accuracy (%)
Figure  1(a) VGG-16 CIFAR-10 3 1,2,4 [2,1,2,1,2,1,1, 1,1,1,4,4,2,4,4] 2.13 93.73
Figure  1(b) ResNet-18 CIFAR-10 2 1,4 [4,4,4,1,4,4,4,1,4,1,1,1,1,1,1,1,4] 3.14 95.34
Figure  1(c) VGG-16 CIFAR-100 3 1,4,8 [8,4,4,1,4,4,1,1 ,1,1,1,4,8,8,8] 4.54 74.07
VGG-16 ImageNet 3 16,8,4 [16,16,16,8,8,8,8,8,4,4,4,4,4,4,4] 9.92 71.188
VGG-16 ImageNet 2 16,8 [16, 16, 16, 8, 16, 8, 8, 8, 8, 8, 8, 8, 8, 8, 8] 12.40 73.98
ResNet-34 ImageNet 2 16,8 [16, 16, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 16, 8, 8, 8, 8, 8, 8] 12.94 75.78
Table 5: SNN accuracy and inference timesteps comparison with QCFS and SEENN-1.
PASC formulation with uniform Quantization Step PASC with Adaptive Layerwise (AL) activation
@@
PASC ($T_{norm}$) 4.21 95.61% SEENN-1 2.01 95.08%
CIFAR-100 QCFS 16 72.80% PASCAL ($T_{eff}$) 3.14 95.34%
PASC ($T_{norm}$) 4.21 76.06% CIFAR-100 SEENN-1 6.19 65.48%
ResNet-34 ImageNet QCFS 64 72.35% PASCAL ($T_{eff}$) 2.63 76.19 %
@@
SEENN-1 29.53 71.84% PASCAL ($T_{eff}$) 4.54 74.07%
PASC ($T_{norm}$) 8.17 74.31% VGG-16 ImageNet PASCAL ($T_{eff})$ 12.40 73.98 %
VGG-16 ImageNet QCFS 1024 74.32% ResNet-34 ImageNet PASCAL ($T_{eff})$ 12.94 75.78 %
PASC ($T_{norm}$) 18.37 74.22%
Table 6: SNN accuracy and inference timesteps comparison with other state-of-the-art works.
Prior Work Dataset Model Accuracy of Prior Work(%) Accuracy of PASC(%)
hao2023bridginggapannssnns CIFAR-100 VGG-16 76.77 (T=16) 76.77 ($T_{norm}=8.79$)
hao2023bridginggapannssnns ImageNet ResNet-34 73.93 (T=32) 74.31 ($T_{norm}=8.17$)
li2023unleashingpotentialspikingneural CIFAR-10 ResNet-18 94.27 (T=11.51) 96.51 ($T_{norm}=8.22$)
li2023unleashingpotentialspikingneural ImageNet VGG-16 73.30 (T=49.54) 74.22 ($T_{norm}=18.37$)
wu2024ftbcforwardtemporalbias ImageNet ResNet34 71.66 (T=16) 74.31 ($T_{norm}=16.67$)
@@ 4.3 Adaptive Layerwise (AL) Activation Results
Figure 1 shows the layerwise AL metric (Equation 6) for different models and datasets. We perform clustering with various values of $\chi$, and compute the quantization step per layer. We also evaluate the effective inference timesteps, $T_{eff}$, which is specified by
$T_{eff}=\frac{1}{|\mathbf{L_{tot}}|}\cdot\sum_{l}r_{E}^{l}\cdot T^{l},$ (16)
where $\mathbf{L_{tot}}$ is the total number of layers and $T^{l}$ is the number of time-steps of the corresponding vanilla SNN for the same layer. Table 4 shows that adapting the quantization step in a layerwise manner achieves competitive accuracy using fewer effective timesteps, thereby improving temporal efficiency. For instance, ResNet-18 provides 95.34% accuracy on CIFAR-10 using $T_{eff}\approx 3.14$, which is comparable to the 96.51% accuracy achieved with a uniform $T_{norm}=8.22$ (see Table 6).
@@ 4.4 Comparison with the State-of-the-Art
We compare our results, presented in Table 5, with those of the QCFS framework (bu2023optimal) and the SEENN-1 method (li2023seenn). In addition to these baselines, we also benchmark our approach against various other state-of-the-art works listed in Table 6. Our results indicate that PASCAL consistently yields SNNs that provide competitive accuracy using fewer timesteps across models and datasets, outperforming the QCFS framework (bu2023optimal). We report comparable performance to the SEENN-1 method (li2023seenn) on CIFAR-10. However, we demonstrate significant performance improvements, both in terms of accuracy and effective timesteps, for CIFAR-100 and ImageNet. This trend is also observed when comparing our method with more recent works in Table 6: PASCAL remains competitive on smaller datasets and achieves superior results on larger ones. The efficacy of PASCAL improves with dataset complexity since it relies on mathematically equivalent ANN-SNN conversion.
@@ 5 Conclusion
In this paper, we proposed PASCAL, which is a precise and temporally efficient ANN-SNN conversion method. We showed that the proposed methodology enables ANNs to be converted into SNNs with a guarantee of mathematical equivalence, leading to near-zero conversion loss. We formally proved that the converted SNN performs the same computations, unrolled over timesteps, as the source ANN with QCFS activation. We demonstrated the energy efficiency of the proposed approach using an analytical model. In addition, we also proposed an Adaptive Layerwise (AL) activation methodology to further improve temporal efficiency during SNN inference. We presented an algorithmic approach to choose an optimal precision for each layer, which reduces the effective number of inference timesteps while providing competitive accuracy.
6 Acknowledgements
This research was supported in part by the RISC-V Knowledge Centre of Excellence (RKCoE), sponsored by the Ministry of Electronics and Information Technology (MeitY). We would like to acknowledge the Robert Bosch Centre for Data Science and AI (RBCDSAI) and the Centre for Development of Advanced Computing (C-DAC) for providing us GPU computing resources. Finally, we would like to acknowledge the work of Varun Manjunath (Electrical Engineering, IIT Madras, Class of 2025) for benchmarking our work on neuromorphic architectural simulators.
@@ A.1 Proof for equivalence between the PASC formulation and an ANN with QCFS activation
For the input layer, the MatMul operation performed in the PASC formulation is the same as the corresponding ANN with QCFS activation. Let the output after the MatMul Operation be $X$. Consider a single element $x\in X$. In the case of ANN, the output after QCFS is
@@
Upon further simplification, we obtain
[…]
Algorithm 2 specifies the Integrate-and-Fire algorithm for the input layer. It also first generates $\widehat{h}(X)=z^{l}$ (refer to Line 3). Subsequently, it generates the final spike train $s^{l}$. We denote the spike train for a given input $x\in X$ by $s^{l}_{x}$. Note that $z^{l}$ only consists of integral multiples of $\theta^{*}=\frac{\theta_{n}}{L_{n}}$. Therefore, after repeated application of the Heaviside step function on $x$, and stacking the output over time, we get
@@
Now, $\sum_{i=1}^{L_{n}}s^{l}_{x}(i)=c\cdot\frac{\theta_{n}}{L_{n}}=\widehat{h}(x)$, $\forall x\in X$. Therefore, $\sum_{i=1}^{L_{n}}s^{l}(i)=z^{l}$, proving part (a). Using Equation 19, we can see that $s_{x}^{l}$ contains values from the set $\{0,\frac{\theta_{n}}{L_{n}}\}$ for all values of $x\in X$. This proves part (b) of the theorem. ∎
Proof for Theorem 3.4. We begin with the remark that $\theta_{n}=\lambda^{l}$ after the training is complete. Consider a single element $x\in z^{l}$ and the corresponding elements $x_{i}\in z^{l}_{i},\forall i\in\{1,2,\cdots L_{n-1}\}$. In an ANN with QCFS activation (see Figure 2),
[…]
Figure 2: A sample QCFS activation layer with $L=4$ and $\theta_{n}=0.25$.
Table 7: Mapping between the value of $s$ at the end of stage 2, and the final number of spikes at the end of stage 3.
@@
$0\geq s\geq-(L_{n}-1)$ 0
After the first stage of Algorithm 1, let the value of $s$ be $k_{1}\cdot\frac{\theta_{n}}{L_{n}}$. If there was a spike in the last timestep, then we obtain an inequality of the form $\frac{\theta_{n}}{2\cdot L_{n}}+\sum_{i}x_{i}-k_{1}\cdot\frac{\theta_{n}}{L_{n}}\geq\frac{\theta_{n}}{L_{n}}$. Upon simplification, we get
$\sum_{i=1}^{L_{n-1}}x_{i}\geq(k_{1}+\frac{1}{2})\cdot\frac{\theta_{n}}{L_{n}}$ (21)
On the other hand, if there was no spike in the last timestep, we instead obtain
$\sum_{i=1}^{L_{n-1}}x_{i}<(k_{1}+\frac{1}{2})\cdot\frac{\theta_{n}}{L_{n}}$ (22)
Although we have a one-directional inequality in either case, we nevertheless need more processing to obtain the second direction of the inequality to choose the correct level. We consider the aforementioned two cases.
Case 1. Let Inequality 21 be the final inequality after stage 1. Inhibitory spikes are not possible in this case. The QCFS function clips the activation if the accumulated input $\sum_{i}x_{i}$ is greater than $(L_{n}-\frac{1}{2})\cdot\frac{\theta_{n}}{L_{n}}$. Therefore, the IF step needs to be done without input at least $T_{1}=L_{n}-1$ times to ensure that the correct level is obtained for all inputs. The bound arises because the minimum value of $k_{1}=0$. At the end of stage 2, let the value of $s$ be $k_{2}\cdot\frac{\theta_{n}}{L_{n}}$. Since an inhibitory spike is not possible, we get
Also, if $k_{2}<L_{n}$, then we also have the following inequality.
Otherwise, we only have the inequality specified in Equation 23. In effect,
Case 2. If Inequality 22 is the final inequality, we need inhibitory spikes to determine the correct level. This is different from Case 1, because we have an upper bound on the value of $\sum_{i=1}^{L_{n-1}}x_{i}$. In effect, any level below $k_{1}$ should be possible. Since $k_{1}\leq L_{n-1}-1$, this stage requires $T_{2}=L_{n-1}-1$ timesteps. Here, normal spikes are not possible. Using a similar analysis as presented in Case 1, we obtain
Stage 3 is used to generate a final spike train from the value of $s$ obtained after stage 2. This layer is used to generate a stacked output $\widetilde{h}(X)$. As a result of this computation, we get that $\sum_{j}\widetilde{h}(X)[j]=min(max(s,0),L_{n})$. The final number of spikes, parametrized by the value of $s$ at the end of stage 2, is given in Table 7. Using 25 and 26, we can conclude that
Proof for Theorem 3.3 We begin by simplifying the right-hand side.
@@
$\sum_{i=1}^{L}\gamma\cdot\frac{z_{i}^{l}*W}{\sqrt{\sigma^{2}+\epsilon}}+\frac{1}{\sqrt{\sigma^{2}+\epsilon}}\sum_{i=1}^{L}\frac{b-\mu}{L}+\sum_{i=1}^{L}\frac{\beta}{L}$ (29)
$S_{1}+S_{2}+S_{3}$ (30)
@@
$\frac{\gamma}{\sqrt{\sigma^{2}+\epsilon}}\cdot z^{l}*W$ (32)
$\frac{\gamma}{\sqrt{\sigma^{2}+\epsilon}}\cdot(z^{l+1}-\beta)\cdot\frac{\sqrt{\sigma^{2}+\epsilon}}{\gamma}-b+\mu$ (33)
$z^{l+1}-\beta-b+\mu$ (34)
@@
Consider an input $z$ to the models $\mathcal{M}$ and $\mathcal{N}$. We invoke Theorems 3.5, 3.4 and 3.3 to show that Invariant 2 holds after the final layer. In other words, after the final layer, $z^{l}=\sum_{i=1}^{L}z_{i}^{l}$, where $L$ is the quantization step of the final layer $l$. In an SNN, the output probabilities are derived by taking the mean over timesteps after the final layer. Therefore, the tensor used for classification in the PASC formulation is $z^{*}=\frac{1}{L}\cdot\sum_{i=1}^{L}z_{i}^{l}=\frac{1}{L}\cdot z^{l}$. Therefore,
@@
$\frac{\frac{z^{l}[c]}{L}}{\frac{\sum_{c}z^{l}[c]}{L}}$ (40)
$\frac{z^{l}[c]}{\sum_{c}z^{l}[c]}$ (41)
$f_{\mathcal{N}}^{z}(c)$ (42)